OpenStudy (anonymous):

Solve: -1/2x + 3 = −x + 7

OpenStudy (australopithecus):

where are you getting stuck

OpenStudy (australopithecus):

fractions aren't hard to deal with just remember this rule 

OpenStudy (australopithecus):

so first things first -1/2x + 3 = −x + 7 = -1/2x + 3(2x)/2x = −x + 7 = (-1+3(2x))/2x = -x + 7 = (-x + 7)2x = -1 +3(2x) = (-x+7)(2x) - 3(2x) = -1

OpenStudy (australopithecus):

you can manipulate fractions by multiplying the numerator and and denominator by the same thing, because it is just like multiplying the fraction by 1

OpenStudy (australopithecus):

by doing that you can easily work with fractions

OpenStudy (australopithecus):

-with that rule there are 3 other rules you need to know, you can combine fractions if they have the same denominator in adding or subtracting (or split them up). -when multiplying fractions just use this rule  -division is the same as multiplication but you flip the bottom fraction before multiplying it